In addition to sine, cosine, and tangent, we have the trigonometric functions secant (sec), cosecant (csc), and cotangent (cot). We define these as follows:
sec x =1/{cos x}
csc x =1/{sin x},
cot x = {cos x}/{sin x},
for those values of x where the right side is defined.
Explain why we must have (cot^2)x + 1 = (csc^2)x for any x such that x is not an integer multiple of 180 degrees.

Answer by Alan3354(69443)

Explain why we must have (cot^2)x + 1 = (csc^2)x for any x such that x is not an integer multiple of 180 degrees.
=================
I can prove it's true, not sure "why we must have it..."
(cot^2)x + 1 = (csc^2)x
cos^2/sin^2 + 1 = 1/sin^2
Multiply by sin^2
cos^2 + sin^2 = 1 --- the Pythagorean Identity.